OpenStudy (anonymous):

simplify. sqrt of... (175x^2y^3)/18 I don't know where to start with this except factors of 175 are 25 and 7, factors of 18 are 9 and 2

OpenStudy (anonymous):

\[\sqrt{(177x ^{2}}y ^{3})/18\] I don't know how, in the equation thing, to display it all over the 18 either

OpenStudy (anonymous):

oops 175...

OpenStudy (anonymous):

\[\sqrt{\frac{175x^{2}y^{3}}{18}}\] \[\sqrt{\frac{25\times7x^{2}y^{3}}{9\times2}} = \frac{5xy}{3}\sqrt{\frac{7y}{2}}\]

OpenStudy (anonymous):

ah okay, just pull it out still in fraction form... how did you display that? I can't figure it out...

OpenStudy (anonymous):

\[\sqrt{175x ^{2}y ^{3}} = \sqrt{5^2*7*x ^{2}*y ^{2}*y} = 5*x*y \sqrt{7y} = 5xy \sqrt{7y}\]

OpenStudy (anonymous):

for fraction the code you need is frac {numerator}{denominator}

OpenStudy (anonymous):

\[\sqrt{18} = \sqrt{3*3*2} = 3\sqrt{2}\]

OpenStudy (anonymous):

ah thanks :) and thanks Harkirat, because I've been practising those since you taught me... the fraction was scaring me... lol

OpenStudy (anonymous):

so u get 5xy --- root(7y/2) 3

OpenStudy (anonymous):

Joz, just practice more.... you will find them fun and easy to do i suggest that u first make factors like I do, then take out pairs (in case of square-root) and then simplify if there r common factors.....

OpenStudy (anonymous):

yeah it makes it easier and it makes it make more sense :)

OpenStudy (anonymous):

\[\sqrt{\frac}\] just testing the display.... \[\frac{175x ^{2}y ^{3}}{18}\]

OpenStudy (anonymous):

\[\sqrt{\frac{175x ^{2}y ^{3}}{18}}\] cool